Shana put $25,000 in an education account on the day her daughter was born. If the account earned 6.75% interest compounded quarterly, what was the total
in the account when her daughter turned 18? Round the final answer to the nearest
cent.

Answer by stanbon(75887)

Shana put $25,000 in an education account on the day her daughter was born. If the account earned 6.75% interest compounded quarterly, what was the total
in the account when her daughter turned 18? Round the final answer to the nearest
---------------------
A(t) = P(1+(r/n))^(nt)
----------
A(18) = 25000(1+(0.0675/4))^(4*18) = $83407.49
--------------------
Cheers,
Stan H.
-------------